OpenStudy (anonymous):

what is the distance between the origin of a Cartesian Coordinate system and the point (-6, 7)?

jimthompson5910 (jim_thompson5910):

I recommend plotting the points (0,0) and (-6,7). Then drawing a right triangle and finding the hypotenuse of the right triangle.

OpenStudy (anonymous):

so i do A2+ B2 = C2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 if that's what you meant, then yes

OpenStudy (anonymous):

yeah so it would then be -6+7= c^2

jimthompson5910 (jim_thompson5910):

what's the length of the blue line segment and green line segment?

OpenStudy (anonymous):

the blue line segment is 7 and the green is -6

jimthompson5910 (jim_thompson5910):

lengths are positive

jimthompson5910 (jim_thompson5910):

so the green is 6 units long

jimthompson5910 (jim_thompson5910):

a = 6 b = 7 c = ??

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 6^2 + 7^2 = c^2 ... ... ... c = ???

OpenStudy (anonymous):

c= 85^2

jimthompson5910 (jim_thompson5910):

actually c^2 = 85 c = ???

OpenStudy (anonymous):

42.5?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

c^2 means c squared NOT c times 2

jimthompson5910 (jim_thompson5910):

what undoes squaring?

OpenStudy (anonymous):

im lost now ._.

jimthompson5910 (jim_thompson5910):

what operation is the complete opposite of squaring

jimthompson5910 (jim_thompson5910):

say you go from 6 to 36 by squaring the 6 how do you go from 36 back to 6?

OpenStudy (anonymous):

square rooting

jimthompson5910 (jim_thompson5910):

so you have to take the square root of both sides of c^2 = 85

jimthompson5910 (jim_thompson5910):